40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Nov 7 17:44:10 2014, diego, Update, LSC, IR Resonance Script Status wavelets.nb.zip
    Reply  Wed Dec 3 01:18:15 2014, diego, Update, LSC, IR Resonance Script Status 
       Reply  Wed Dec 3 02:01:57 2014, Koji, Update, LSC, IR Resonance Script Status 
Message ID: 10747     Entry time: Wed Dec 3 01:18:15 2014     In reply to: 10687     Reply to this: 10749
Author: diego 
Type: Update 
Category: LSC 
Subject: IR Resonance Script Status 

Tonight I started testing a new method for the fine scan:

  • the idea is to use the zero crossings of the PO*11_ERR_DQ signals after (or as an alternative of) the fine scan, but such signals are quite dirty, so I need to find some good way to smooth/filter them;
  • I didn't manage to make many tests, because:
    • once arms were locked fine with ALS, the CARM & DARM lock wasn't very robust, in both acquiring and maintaining lock;
    • during the night, the slow OFSs of the arms misbehaved, and at least once per arm they raised their warning box (independently from each other, and it was hastily recovered), even for values that had been perfectly fine before; I am confused about this;
    • as a result, notwithstanding many tries, the beatnotes are gone;
  • I have enough information to push the script a little further, but I'll do more testing soon;

 

ELOG V3.1.3-